OpenStudy (anonymous):

given f(x)=x^2=8 and g(x)=x-1/x find (g of f)(-4)

OpenStudy (anonymous):

f(x)=x.2 or x^2=8 :s

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

joe something looks funny about your question

OpenStudy (anonymous):

joseph*

OpenStudy (anonymous):

g(f(x))=(x^2-1)/x^2 ((-4)^2-1)/(-4)^2 15/16

OpenStudy (anonymous):

my bad *f(x)=x^2+8

OpenStudy (anonymous):

too bad, i did wrong question :p

OpenStudy (anonymous):

(x^2+(8-1))/((X^2)+8) (16+7)/(16+8) 23/24